### v1

# From Stimuli to Minds: Enhancing Psychological Reasoning in LLMs via Bilateral Reinforcement Learning

Abstract
Large Language Models show promise in emotion understanding, social reasoning, and empathy, yet they struggle with psychologically grounded tasks that require inferring implicit mental states in context-rich, ambiguous settings. These limitations arise from the absence of theory-aligned supervision and the difficulty of capturing nuanced mental processes in real-world narratives. To address this gap, we leverage expert-labeled, psychologically rich scenarios and propose a trajectory-aware reinforcement learning framework that explicitly imitates expert psychological thought patterns. By integrating real-world stimuli with structured reasoning guidance, our approach enables compact models to internalize social-cognitive principles, perform nuanced psychological inference, and support continual self-improvement. Comprehensive experiments across multiple benchmarks further demonstrate that our models achieve expert-level interpretive capabilities, exhibiting strong out-of-distribution generalization and robust continual learning across diverse, challenging, and psychologically grounded tasks.
1 Introduction
Large Language Models (LLMs) show strong generalization across diverse language tasks (Wu 2025; Zhang et al. 2025; Feng et al. 2025). Their emerging potential in psychological domains, such as emotion understanding (Kovacevic et al. 2024), social reasoning (Leng and Yuan 2023), and empathy recognition (Sorin et al. 2024) has drawn growing research interest. These tasks require not only linguistic comprehension but also nuanced inference of implicit mental states and emotional cues, often without explicit supervision or clearly defined ground truths. Unlike conventional language tasks, they involve rich psychological stimuli embedded in ambiguous, socially grounded scenarios shaped by diverse cultural norms, interpersonal dynamics, and lived experiences. Although LLMs show partial sensitivity to such cues, performance on psychologically grounded tasks remains limited from human-level competence (Ke et al. 2025).
Figure 1: A sample ToMbench question presents two psychological stimuli—one focused on direct emotion prediction, the other requiring nuanced reasoning about mental states.
Figure 2: The figure summarizes key psychological stimuli across our datasets—emotional variables, narrative transformations, and collective factors—comprising 35,084, 24,831, and 36,890 QA instances, each offering insights into human narratives.
To evaluate LLMs’ capacity for cognitive reasoning, several benchmarks have emerged. ToMbench provides a structured multiple-choice framework for assessing Theory of Mind (ToM) across eight task types and 31 social-cognitive capabilities (Chen et al. 2024), and shows that even advanced models like GPT-4o (Achiam et al. 2023) still lag behind human performance. Other benchmarks such as Psychobench (Li et al. 2024) and CogBench (Coda-Forno et al. 2024) further explore emotion inference and pragmatic reasoning. These evaluations mark a growing effort to systematize measurement of social-cognitive skills in LLMs. While LLMs display partial sensitivity to mental states, their reasoning remains fragmented, often relying on surface cues rather than robust mental models and struggle to generalize in complex scenarios, falling short of human-like social understanding.
Despite recent advances, three key challenges remain. (1) High quality data scarcity: Benchmarks like ToMbench are small in scale (Wu et al. 2023), and many datasets use LLM-generated content (Hu et al. 2024), limiting their value for tuning psychological reasoning (Long et al. 2024). (2) Reasoning mismatch across tasks: Psychological tasks differ in cognitive demands. Theories as Epstein’s (Epstein 1998) and Fuzzy-Trace Theory (Reyna and Brainerd 1998) separate fast, intuitive reasoning from slow, analytical thought. Uniform strategies may hurt performance on intuition-driven tasks (Ji et al. 2025). As an example illustrated in Figure 1. (3) Poor generalization in small models: Compact LLMs tend to overfit and struggle to generalize. Many depend on LLM-generated labels or costly prompting (Wang et al. 2024).
To address these limitations, we make three contributions. First, we construct a large-scale dataset, StimuliQA, grounded in professional psychological theory and real-world interviews. It features over 3,000 annotated stimuli with 58 psychological variables, which are subsequently converted into question–answer (QA) pairs, as detailed in Figure 2. Second, we introduce Psy-Interpreter, a reinforcement learning (RL) framework inspired by dual-system psychological theories. It incorporates a trajectory cache and bilateral reasoning (BR) to foster expert-like psychological analysis across diverse tasks. Third, we demonstrate that compact models, when trained with the Psy-Interpreter framework and rationale-augmented supervision, can rival much larger systems across multiple benchmarks, highlighting the effectiveness of structured RL and reasoning-aware supervision in enhancing efficiency and generalization for LLMs.
We validate our approach through three experiments. First, our expert-labeled dataset consistently improves the out-of-distribution (OOD) performance of mainstream post-training methods, clearly surpassing existing datasets without expert annotations. Second, we show that the Psy-Interpreter framework with dual-system training substantially outperforms standard generation or training methods across multiple diverse task types. Third, we show that models trained under our framework not only generalize better to unseen psychological tasks, but also possess self-annotation capabilities that enable continual learning, thereby closing the gap between expert supervision and autonomous psychological reasoning.
Together, these contributions form a unified and practical framework for advancing social-cognitive reasoning in LLMs through real-world data, psychologically informed learning, and efficient model optimization. Based on our experiments, we argue that explicit knowledge injection and the imitation of expert psychological thought patterns collectively endow the model with expert-level interpretive capabilities.
2 Related Works
2.1 LLMs in Psychological Tasks
LLMs are increasingly evaluated on psychological tasks involving theory of mind and moral reasoning. Benchmarks show that even top models such as GPT-4o underperform humans in belief reasoning, indicating reliance on superficial cues (Xiao et al. 2025; Shapira et al. 2024). Datasets such as SimpleToM, CogBench, and SocialIQa extend evaluations to emotion and social inference (Gu et al. 2024; Coda-Forno et al. 2024; Sap et al. 2019b). In moral reasoning, LLMs may align with or even surpass humans in perceived ethical competence (Dillion et al. 2025; Huang et al. 2023), though often via pattern matching. Empathy studies find LLMs capable of emotionally appropriate responses, occasionally preferred over human ones (Ayers et al. 2023; Sorin et al. 2024), albeit with limited contextual nuance (Yang, Ye, and Du 2024). Open models still lag behind proprietary ones (Li et al. 2024), highlighting the need for theory-informed benchmarks and training methods (Xie et al. 2024; Qiu et al. 2024).
Figure 3: The framework comprises: StimuliQA, stimuli with expert psychological labels; Psy-Interpreter, a training framework tailored for psychological tasks; and Continual Learning, demonstrating continual learning capability through self-labeling.
2.2 CoT and Reinforcement Learning
Chain-of-Thought (CoT) prompting improves deliberate reasoning by encouraging stepwise thinking (Wei et al. 2022; Kojima et al. 2022), and is often combined with RL to align outputs with human preferences (Ouyang et al. 2022). Process supervision and gradient-level feedback, such as in DeepSeek-R1, further enhance intermediate reasoning quality (Liu and et al. 2025). Frameworks like Reflexion and Tree of Thoughts introduce self-evaluation and planning to strengthen coherence (Shinn et al. 2023; Yao et al. 2023). Despite progress, challenges remain in balancing fluency with rigor and designing scalable feedback for ambiguous reasoning. CoT and RL provide a strong basis for aligning LLMs in psychological reasoning tasks and broader real-world applications.
3 Methodology
This section outlines our methodology for enhancing psychological reasoning as shown in the Figure 3. We first construct StimuliQA, a dataset pairing human-labeled narratives with LLM-generated QA. We then propose Bilateral Reinforcement Learning, integrating token accuracy, format compliance, reasoning depth, and repetition control into a unified reward. A Trajectory Cache stabilizes training via recent rollout tracking and bonus conditions. Finally, Continuous Learning enables refinement through confident predictions.
3.1 Data Construction
We built StimuliQA to support psychological reasoning by combining expert annotations, aiming to inject expert knowledge and guide LLMs toward human-like reasoning. Comparison with other datasets can be found in the Appendix.
Stimuli Collection and Annotation
We collected 3,280 real-life stimuli with emotional and social content. Human annotators labeled them based on variables above, and utilized LLMs to generate QA pairs from the labels, ensuring expert grounding and generative diversity.
Psychological Variable Design
The dataset comprises 58 variables across three dimensions: Emotional Reactions (29 variables), which represent affective and social-emotional responses grounded in Lazarus’s appraisal theory (Lazarus 1991) , and further explored in settings as (Han 2025); Narrative Transformation (12 variables), reflecting tone shifts and redemptive arcs, inspired by McAdams et al. (McAdams et al. 2001); and Collective Psychology (17 variables), covering indicators of self-worth and community connection based on Ryff & Keyes’s model of psychological well-being (Ryff and Keyes 1995). This variable structure enables exploration of how individuals process and narrate life experiences.
3.2 Bilateral Reinforcement Learning
We propose a Bilateral RL frame with a Trajectory Cache that adopt Trajectory-aware GRPO based on Group Relative Policy Optimization (GRPO) (Shao et al. 2024) and Bilateral Reward stabilizes rewards and promotes structured reasoning.
Trajectory Cache
Fixed rewards are often suboptimal across model scales or tasks. We use a Trajectory Cache to track recent performance and adjust rewards by trend, stabilizing estimation. With $B$ batches and cache size $C$, the batch caches $B_{c}=B\times C$ summarize training dynamics.
T-GRPO Objective
The objective of T-GRPO (Trajectory-aware Group Relative Policy Optimization ) is defined as:
$\!\mathcal{J}_{\text{T-GRPO}}(\theta)\!=\!\frac{1}{B_{c}G}\sum_{b=1}^{B_{c}}\sum_{i=1}^{G}\frac{1}{|o_{b,i}|}\!\sum_{t=1}^{|o_{b,i}|}\!\bigg{[}\min\bigg{(}r_{b,i,t}(\theta)\hat{A}_{b,i,t},$ (1)
$\qquad\text{clip}(r_{b,i,t}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{b,i,t}\bigg{)}\bigg{]}\!-\!\beta D_{\text{KL}}(\pi_{\theta}\,\|\,\pi_{\text{ref}}),$
where the importance ratio and normalized advantage are:
$r_{b,i,t}(\theta)=\frac{\pi_{\theta}(o_{b,i,t}\mid q_{b},o_{b,i,<t})}{\pi_{\theta_{\text{old}}}(o_{b,i,t}\mid q_{b},o_{b,i,<t})},\ \hat{A}_{b,i,t}=\frac{r^{\text{Final}}_{b,i}-\bar{r}}{\sigma_{r}+\epsilon},$ (2)
with $\bar{r}$ and $\sigma_{r}$ being the batch-level mean and standard deviation computed over all group-level final rewards $r^{\text{Final}}_{b,i}$ as defined in the Bilateral Reward section.
Bilateral Reward
For our overall bilateral reward, each group $(b,i)$ is assigned a final reward for $r^{\text{Final}}_{b,i}$ listed here:
$r^{\text{Final}}_{b,i}=w^{\text{F1}}\cdot r^{\text{F1}}_{b,i}+w^{\text{fmt}}\cdot r^{\text{fmt}}_{b,i}+r^{\text{BR}}_{b,i}-r^{\text{rep}}_{b,i},$ (3)
where $w^{(\cdot)}$ denotes the weight and $r^{(\cdot)}_{b,i}$ denotes the reward term, which will be introduced in further detail below.
Answer Quality ($r^{\text{F1}}$)
We compute $r^{\text{F1}}_{b,i}$ as the token-level F1 score between the predicted answer $\text{Ans}^{\text{pred}}_{b,i}$ the corresponding ground-truth answer $\text{Ans}^{\text{gold}}_{b,i}$. Formally:
$r^{\text{F1}}_{b,i}=\frac{2P_{b,i}R_{b,i}}{P_{b,i}+R_{b,i}},$ (4)
$\ P_{b,i}=\frac{|y^{\text{pred}}_{b,i}\cap y^{\text{gold}}_{b,i}|}{|y^{\text{pred}}_{b,i}|},\ R_{b,i}=\frac{|y^{\text{pred}}_{b,i}\cap y^{\text{gold}}_{b,i}|}{|y^{\text{gold}}_{b,i}|},$ (5)
where $y^{\text{pred}}_{b,i}$ $y^{\text{gold}}_{b,i}$ are token sets after normalization, including lowercase conversion, punctuation and space cleanup.
Format Compliance ($r^{\text{fmt}}$)
To ensure structural consistency, we define a binary reward for format correctness:
$r^{\text{fmt}}_{b,i}=\begin{cases}1,&\text{if }o_{b,i}\text{ matches predefined format constraints,}\\ 0,&\text{otherwise.}\end{cases}$ (6)
Format validation checks if the response is in the correct format ( <think>… </think> <answer>… </answer> ).
Bilateral Reasoning ($r^{\text{BR}}$)
[…]
$\bar{L}=\frac{1}{B_{c}G}\sum_{b=1}^{B_{c}}\sum_{i=1}^{G}L_{b,i},\quad\bar{r}^{\text{F1}}=\frac{1}{B_{c}G}\sum_{b=1}^{B_{c}}\sum_{i=1}^{G}r^{\text{F1}}_{b,i}.$ (7)
We then compute the bilateral reward as:
$r^{\text{BR}}_{b,i}=\begin{cases}\delta,&\frac{L_{b,i}}{\bar{L}}<\tau_{-}\ \text{and}\ r^{\text{F1}}_{b,i}>\bar{r}^{\text{F1}},\\[2.0pt] \delta,&\frac{L_{b,i}}{\bar{L}}>\tau_{+}\ \text{and}\ r^{\text{F1}}_{b,i}>\bar{r}^{\text{F1}},\\[2.0pt] 0,&\text{otherwise,}\end{cases}$ (8)
where $\tau_{-}$ and $\tau_{+}$ denote the lower and upper thresholds, respectively. This reward encourages the model to adopt an appropriate reasoning length, using concise responses for simpler cases and extended reasoning for complex ones.
Further, comparing $r^{\text{F1}}_{b,i}$ with $\bar{r}^{\text{F1}}$ ensures that length-based rewards are granted only to outputs that are both accurate and reliably reasoned, preventing the model from receiving incentives for merely producing longer or shorter responses without demonstrating genuine understanding.
Repetition Penalty ($r^{\text{rep}}$)
To discourage verbosity and repetition, we compute a penalty over the <think> segment based on the ratio of repeated to total 4-grams:
$r^{\text{rep}}_{b,i}=-\min\left(\tau_{\text{rep}},\frac{\text{Repeat}_{4}(o^{\text{think}}_{b,i})}{\text{Total}_{4}(o^{\text{think}}_{b,i})+\epsilon}\right),$ (9)
where $\tau_{\text{rep}}$ denotes the maximum penalty score for penalizes semantic redundancy and promotes meaningful reasoning.
Other Rewards
Additional general reward formulations-including the base reward, short reward, length reward and the length-based reward with repetition penalty (RP) were also utilized and systematically compared in the later section (see Section 4). Full details are provided in the Appendix.
3.3 Continuous Learning
To enable self-evaluation and self-improvement under low-resource settings, we propose a simple framework for continuous learning. The model not only generates predictions but also assesses their quality and improves over time—producing more user-aligned and psychologically appropriate responses. We define the learning criterion as:
[…]
where $x$ is a model output, $\text{valid}(x)$ verifies format and $\text{confidence}(x)$ is the model’s estimated confidence score.
4 Experiments
This section is organized around five core research questions (RQs): RQ1: Can human-labeled psychological data improve model generalization under various training methods? RQ2: Do tailored reward functions enhance recognition of explicit and latent psychological states? RQ3: Does Psy-Interpreter perform well on OOD datasets? RQ4: Is continual learning generally effective? RQ5: Can Psy-Interpreter reliably identify if a psychological question requires reasoning?
4.1 Experimental Setup
Datasets
We used six different datasets: our StimuliQA and five OOD sets—ToMbench (Chen et al. 2024), SimpleToM (Gu et al. 2024), SocialIQa (Sap et al. 2019a), CosmosQA (Huang et al. 2019), and selected BIG-Bench Hard (Suzgun et al. 2023) subsets (disambiguation_qa, formal_fallacies, causal_judgement). For each, we extracted the correct option to construct QA pairs grounded in narrative context. As SimpleToM and SocialIQa answers were difficult to connect to their questions, we appended them to the question text. We sampled 200 training and 100 test examples per psychological parameter, excluding four low-frequency ones (C_acol, C_agov, C_anet, C_apol), yielding 10,800 training and 5,400 test instances left for training and testing proposes.
Metrics
Consistent with prior QA research like(Deutsch, Bedrax-Weiss, and Roth 2021) and (Su et al. 2019), we use F1 score to evaluate our dataset. For OOD benchmarks, we adopt F1-based accuracy: each output is matched to the option with highest F1. If multiple options tie, we return E (“Not Applicable”); accuracy is computed on these selections.
Implementation Details
We conduct three experiments. First, we train models on our dataset using two training paradigms (Table 1, Figure 4) and compare their performance against two constructed comparison datasets, LlamaQA and MistralQA. Second, we train Qwen models on StimuliQA with varied reward functions and decoding strategies, evaluated on the held-out test set (Table 2). Third, we label logical reasoning samples on StimuliQA and use them to SFT-train Qwen models, thereby injecting knowledge without changing the generation format (Mecklenburg et al. 2024). As Reinforcement Learning from Human Feedback (RLHF) primarily re-ranks existing knowledge (Christiano et al. 2017; Yue et al. 2025; Li et al. 2025), we apply our RL framework to the SFT model to produce Psy-Interpreter, evaluated on five OOD datasets and further tested with a continual learning module. Full details are provided in the Appendix.
Training Sets ToMbench SimpleTom SocialIQa CosmosQA BIG-Bench Hard
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Qwen2.5-0.5B-Instruct
LlamaQA 11.95 26.89% 15.37 32.03% 15.55 30.60% 12.29 28.21% 11.06 40.47%
MistralQA 10.01 21.33% 20.43 26.88% 22.79 32.70% 10.30 24.29% 14.45 39.74%
StimuliQA 16.61 34.13% 22.32 53.59% 25.37 40.74% 12.82 30.08% 16.23 48.91%
Qwen2.5-1.5B-Instruct
LlamaQA 12.14 27.52% 20.03 34.93% 21.80 40.28% 12.43 28.74% 12.39 44.10%
MistralQA 10.27 22.62% 25.02 33.74% 27.31 37.87% 11.31 26.33% 14.50 39.88%
StimuliQA 18.26 37.94% 36.21 49.84% 27.32 41.91% 14.23 30.92% 23.51 56.04%
Qwen2.5-3B-Instruct
LlamaQA 15.91 32.69% 18.48 26.16% 34.62 49.59% 15.11 31.96% 13.63 44.98%
MistralQA 14.88 30.70% 35.02 33.33% 35.28 40.84% 12.69 27.34% 19.33 41.05%
StimuliQA 22.06 42.38% 37.62 56.44% 43.40 53.74% 15.99 33.17% 19.74 54.88%
Table 1: Comparison of SFT training on StimuliQA and two other datasets by both accuracy and F1 score.
Figure 4: Comparison of GRPO training on StimuliQA and two other training datasets on Qwen0.5B and 3B. Full table with Qwen1.5B’s results is located in the Appendix.
Model
Specific
Emotions
Targeted
Emotions
Direction of
Change
Self
Identification
Needs
Fulfilment
National
Theme
Overall
Qwen2.5-0.5B-Instruct
Naïve 10.72 9.20 14.60 13.78 10.58 13.76 11.46
CoT 10.95(+0.23) 10.43(+1.23) 14.51(-0.09) 15.30(+1.52) 11.42(+0.84) 14.11(+0.35) 12.31(+0.85)
SFT 14.08(+3.36) 19.70(+10.5) 29.27(+14.67) 21.29(+7.51) 22.07(+11.49) 29.79(+16.03) 21.70(+10.24)
Basic R1 25.25 27.11 27.91 32.78 32.59 30.62 29.23
Length Reward 25.49(+0.24) 26.83(-0.28) 28.40(+0.49) 32.80(+0.02) 31.86(-0.73) 27.81(-2.81) 28.63(-0.60)
Short Reward 22.61(-2.64) 28.90(+1.79) 27.82(-0.09) 34.52(+1.74) 33.80(+1.21) 28.50(-2.12) 29.49(+0.26)
Length w RP 26.70(+1.45) 28.49(+1.38) 29.85(+1.94) 34.59(+1.81) 33.67(+1.08) 30.18(-0.44) 30.37(+1.14)
BR 29.02(+3.77) 31.20(+4.09) 38.70(+10.79) 37.84(+5.06) 39.97(+7.38) 38.74(+8.12) 34.97(+5.74)
Qwen2.5-1.5B-Instruct
Naïve 11.88 10.68 18.06 15.64 13.13 16.67 13.46
CoT 14.87(+2.99) 15.07(+4.39) 22.59(+4.53) 19.09(+3.45) 17.68(+4.55) 20.18(+3.51) 17.37(+3.91)
SFT 19.13(+7.25) 21.48(+10.8) 33.00(+14.94) 31.32(+15.68) 29.43(+16.30) 31.58(+14.91) 26.27(+12.81)
Basic R1 34.29 30.57 37.03 35.28 38.22 36.81 34.46
Length Reward 33.26(-1.03) 30.91(+0.34) 35.94(-1.09) 35.89(+0.61) 38.00(-0.22) 36.75(-0.06) 34.41(-0.05)
Short Reward 34.81(+0.52) 30.56(-0.01) 36.92(-0.11) 37.00(+1.72) 38.78(+0.56) 38.23(+1.42) 35.14(+0.68)
Length w RP 34.90(+0.61) 32.38(+1.81) 37.05(+0.02) 38.01(+2.73) 40.54(+2.32) 39.25(+2.44) 36.32(+1.86)
BR 38.45(+4.16) 36.00(+5.43) 41.95(+4.92) 42.09(+6.81) 43.88(+5.66) 42.52(+5.71) 39.98(+5.52)
Qwen2.5-3B-Instruct
Naïve 13.44 13.27 19.97 16.20 15.19 18.68 15.34
CoT 13.86(+0.42) 15.11(+1.84) 21.60(+1.63) 17.65(+1.45) 16.27(+1.08) 20.03(+1.35) 16.70(+1.36)
SFT 22.45(+9.01) 22.04(+8.77) 34.47(+14.50) 31.68(+15.48) 27.24(+10.97) 32.73(+14.05) 26.95(+11.61)
Basic R1 34.17 31.23 36.81 36.04 37.98 38.71 35.04
Length Reward 33.02(-1.15) 31.56(+0.33) 37.90(+1.09) 36.77(+0.73) 38.78(+0.80) 37.86(-0.85) 35.14(+0.10)
Short Reward 35.46(+1.29) 31.28(+0.05) 38.02(+1.21) 37.84(+1.80) 40.38(+2.40) 37.72(-0.99) 35.81(+0.77)
Length w RP 35.50(+1.33) 33.56(+2.33) 39.88(+3.07) 39.02(+2.98) 41.79(+3.81) 39.80(+1.09) 37.39(+2.35)
BR 39.17(+5.00) 36.25(+5.02) 41.87(+5.06) 42.03(+5.99) 44.44(+6.46) 43.38(+4.67) 40.38(+5.34)
Table 2: Comparison of different training and generation methods on our StimuliQA dataset across six psychological dimensions from our StimuliQA. Our BR reward consistently achieves the best overall performance across all six domains.
Models ToMbench SimpleTom SocialIQa CosmosQA BIG-Bench Hard
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
GPT-4 nano 22.11 46.22% 67.03 54.34% 57.03 68.03% 15.82 35.00% 9.30 41.75%
Claude 3 Haiku 18.81 49.65% 19.32 37.31% 15.94 52.64% 11.28 33.27% 7.07 40.15%
DeepSeek Reasoner 26.63 56.02% 78.98 69.17% 63.93 70.36% 13.13 35.23% 40.52 69.93%
Llama 3.3 70B Instruct 27.42 52.24% 76.43 61.46% 60.91 69.07% 15.27 37.19% 17.61 69.34%
Mistral 8$\times$7B Instruct 12.03 34.46% 12.06 42.52% 12.14 42.23% 7.49 24.66% 5.39 26.42%
Qwen2.5-0.5B-Instruct
Simple RL Zoo 13.38 14.42% 12.89 6.63% 16.56 12.23% 10.33 5.22% 9.12 6.86%
NuExtract 12.69 7.21% 12.20 1.39% 14.44 2.54% 8.02 2.50% 6.83 2.04%
New Merges Serialization 15.21 23.27% 17.30 13.86% 15.14 22.69% 9.80 22.69% 8.04 6.13%
Psy-Interpreter 25.27 40.66% 37.63 65.10% 43.01 51.30% 23.58 38.69% 23.05 48.76%
Psy-Interpreter-SFT 62.28 58.82% 85.25 65.45% 76.58 74.04% 34.43 43.45% 62.41 56.06%
Qwen2.5-1.5B-Instruct
Simple RL Zoo 3.36 4.51% 9.02 6.51% 4.33 7.31% 2.91 2.50% 3.98 8.03%
Nemotron Reasoning 2.02 4.65% 6.38 6.10% 7.30 10.21% 1.69 5.26% 6.98 9.78%
DeepSeek-R1 Distilled 11.39 22.64% 26.80 26.68% 9.25 22.49% 7.33 12.14% 2.73 13.28%
Psy-Interpreter 26.54 49.30% 39.02 59.31% 56.40 63.94% 23.61 41.34% 22.33 50.07%
Psy-Interpreter-SFT 65.94 62.25% 84.95 66.26% 80.55 78.86% 36.37 45.68% 64.04 58.69%
Qwen2.5-3B-Instruct
Transformer RL 16.25 18.86% 15.55 9.56% 25.20 39.33% 11.14 6.53% 9.33 5.26%
Tulu Trained 21.54 22.67% 33.83 16.97% 46.24 39.12% 14.44 24.28% 34.92 29.34%
Raspberry 15.46 15.96% 33.92 24.35% 28.93 26.06% 10.36 3.65% 13.23 6.57%
Psy-Interpreter 28.17 51.33% 56.83 58.73% 65.54 71.09% 25.60 44.33% 18.87 56.06%
Psy-Interpreter-SFT 66.33 58.36% 84.98 67.92% 82.82 81.24% 36.88 47.56% 65.86 60.88%
Table 3: Performance of our Psy-Interpreter and continual learning with SFT on all the OOD datasets, compared against strong commercial models and open-source baselines, with consistent gains across all five datasets.
4.2 RQ1: Effect of Human-Labeled Data
One of the central hypotheses in our work is that human-labeled data, particularly when annotated by trained psychology students, can inject high-quality domain knowledge into compact models during SFT. We evaluate this hypothesis by comparing models trained on our StimuliQA dataset against those trained on synthetic datasets generated by two powerful large models: Llama 3.3-70B and Mistral 8$\times$7B. All datasets were used to train Qwen2.5-Instruct models of three different sizes (0.5B, 1.5B, 3B) via SFT, and evaluated on five diverse OOD benchmarks: ToMbench, SimpleToM, SocialIQa, CosmosQA, and BIG-Bench Hard. Where we have the final results displayed in Table 1 and Figure 4. This comparison reveals performance gaps and tests the transferability of psychological reasoning across model sizes and tasks. These benchmarks span theory of mind, social commonsense, and logical reasoning, allowing for a comprehensive assessment of domain generalization and model capabilities.
As shown in Table 1, models trained on StimuliQA consistently outperform those trained on synthetic data across all model sizes and benchmarks, with especially strong gains in the 3B setting. For instance, on SimpleToM, Qwen2.5-3B achieves 37.62 F1 and 56.44% accuracy, far exceeding Llama 3.3 (18.48/26.16%) and Mistral 8$\times$7B (35.02/33.33%). These substantial margins further reinforce our claim that: human-labeled data conveys more nuanced and structured psychological knowledge than synthetic outputs. The consistent OOD improvements indicate this knowledge is indeed generalizable, highlighting the professional quality and domain relevance of our dataset. Annotations by trained psychology students provide valuable domain expertise essential for effective model training. When used in GRPO training (Figure 4), human-labeled data still offers an advantage.
4.3 RQ2: Impact of Reward Design
Our results in Table 2 show that psychologically grounded reward functions substantially enhance model performance, and Figure 5 also shows a significant training accuracy boost. In particular, the BR reward, which integrates structure, emotion sensitivity, and length normalization, consistently outperforms baselines like Basic R1 or single-aspect rewards. For instance, BR raises overall F1 on Qwen2.5-1.5B from 34.46 to 39.98, with +5.66 and +5.71 F1 gains in Needs Fulfilment and National Theme, respectively. These gains suggest BR promotes not only token-level accuracy but also deeper emotional and moral reasoning. Beyond absolute improvements, BR yields more balanced results across all psychological dimensions, unlike length-based or repetition-penalized rewards, which often improve certain aspects at the cost of degrading others. This highlights BR’s ability to align model behavior with complex, multi-faceted psychological goals.
Moreover, BR scales well across model sizes, offering +5.34 F1 on Qwen-3B and +5.74 on Qwen-0.5B, making it suitable for compute-constrained settings. Finally, BR-trained models generate more coherent moral reasoning, richer self-reflection, and clearer emotional distinctions indicating a shift toward more theory-aligned internal representations.
4.4 RQ3: Generalization of Psy-Interpreter
As shown in Table 3, Psy-Interpreter demonstrates strong zero-shot generalization across five diverse OOD benchmarks after injecting knowledge via SFT and GRPO training with our StimuliQA dataset. Despite being trained without any direct supervision on these test sets, it consistently outperforms other baselines such as Simple RL Zoo, NuExtract, and DeepSeek-R1 Distilled across all model sizes. Notably, the 3B variant achieves 28.17 F1 on ToMbench, 56.83 on SimpleToM, and 65.54 on SocialIQa, marking substantial gains over both in-house and open-source compact models. These results suggest that Psy-Interpreter successfully internalizes psychologically grounded knowledge during training, enabling it to reason beyond its original domain.
Figure 5: Base and Bilateral reward Training comparison.
4.5 RQ4: Effectiveness of Continual Learning
Table 3 shows that our continual learning framework (Psy-Interpreter-SFT) achieves substantial and consistent improvements across all five OOD psychological benchmarks. For the 0.5B model, accuracy rises from 40.66% to 58.82% on ToMbench, 51.30% to 74.04% on SocialIQa, and 38.69% to 43.45% on CosmosQA. F1 scores increase even more sharply, e.g., 25.27 to 62.28 on ToMbench and 23.05 to 62.41 on BIG-Bench Hard, indicating better correctness and closer alignment with human-preferred reasoning. The 1.5B and 3B variants maintain these gains, reaching up to 82.82 F1 on SocialIQa and consistently surpassing their non-SFT counterparts. Despite their compact size, our models rival or exceed larger commercial LLMs: Psy-Interpreter-SFT (3B) outperforms GPT-4 nano (57.03) and Claude 3 Haiku (15.94), and matches or surpasses DeepSeek Reasoner on most tasks.
Despite their smaller scale, our models often remarkably match or notably surpass larger commercial LLMs. Psy-Interpreter-SFT (3B) achieves an F1 of 82.82 on SocialIQa, exceeding GPT-4 nano (57.03) and Claude 3 Haiku (15.94), and outperforming DeepSeek Reasoner on most tasks.
4.6 RQ5: Length Distribution Comparison
Figure 6 illustrates the reasoning length distributions under the bilateral and base reward settings on the Unexpected Outcome Test from ToMbench, which comprises 300 samples with an intuitive question-difficulty hierarchy. We only include samples with non-zero F1 scores, as these responses provide reasoning chains that are more representative and interpretable than those from entirely incorrect outputs. Index 1 corresponds to simple stories with simple questions, Index 2 to the same stories with harder questions, and Index 3 to more complex stories with direct questions. The figure also reports the total number of valid samples under each reward configuration. As shown, the bilateral reward yields a more structured and interpretable distribution: Index 1 responses remain concise, while Index 2 and Index 3 exhibit longer and more differentiated reasoning chains. In contrast, the base reward produces a noisier distribution with reduced separation between simple and complex QA pairs, occasionally generating overly long reasoning even for direct emotion predictions. This indicates that our bilateral reward better aligns reasoning length with task difficulty, enhancing overall interpretability and efficiency. Moreover, the consistent accuracy gains in Tables 2 and 3 underscore the importance of carefully controlling reasoning length in complex psychological reasoning. In addition, our reward method was able to generate more correct samples at each difficulty level, yielding 12 additional samples for simple emotion prediction and 19 and 8 more for the harder ones, respectively, validating its effectiveness.
Figure 1 shows a reasoning samples with BR. More detailed comparisons and analysis can be found in the Appendix where we argued that our training framework successfully improved Emotional Quotient (EQ) of LLMs.
Figure 6: The figure shows the distribution of reasoning lengths under the bilateral reward and the base reward.
5 Conclusion
We first introduce the expert-annotated dataset StimuliQA and Psy-Interpreter, an RL framework firmly grounded in psychological theory that enhances emotional and cognitive reasoning in LLMs. Leveraging high-quality human-labeled data with structured RL substantially improves social-cognitive reasoning in compact models, yielding consistent gains across five diverse OOD benchmarks. Our Bilateral Reward aligns reasoning length with task complexity and improves accuracy, while Psy-Interpreter-SFT with continual learning module effectively narrows the gap with larger commercial LLMs.
Appendix A Training Parameters
For all our experiments, we used four NVIDIA RTX A6000 with memory usage of 49140MiB. And driver version of 550.144.03 and CUDA version of 12.4.
A.1 Prompts for Testing
CoT-related Prompt.
[content | trim] A conversation between User and Assistant. The user provides a “Stimuli:”, and the Assistant reads it and answers the “Question:”. The Assistant first thinks about the reasoning process internally, then provides the answer to the user. The reasoning and answer are enclosed in the following tags:
  • Reasoning: <think> … </think>
  • Answer: <answer> … </answer>
Example: <think> reasoning process here </think> <answer> answer here </answer>
Direct Generation Prompt.
[content | trim] A conversation between User and Assistant. The user provides a “Stimuli:”, and the Assistant reads it and answers the “Question:”. The Assistant only provides the answer, enclosed as:
  • <answer> answer here </answer>
Topic Generation Prompt.
Refer to the following topics as conceptual anchors. Generate as many psychologically or contextually related topics as possible while preserving thematic coherence:
["sample topics"]
Dataset Generation Prompt.
You are an AI assistant that generates one question–answer pair based on a short Stimuli. The question should assess the human-labeled psychological parameter [column_name] ([description]), with value range [range]. Ensure that the question:
  • Directly relates to the Stimuli,
  • Evaluates the concept of [category],
  • Follows the exact format:
Question: <Insert question here>
Answer: <Insert answer here>
Reason: <Explain why the answer is correct>
Language should be clear, concise, and include plausible distractor options.
Comparison Dataset Generation Prompt.
Generate a Stimuli based on ([topic]), then generate the Q&A pair for psychological parameter [column_name] ([description]) in the same format as above.
Reasoning CoT Generation Prompt.
Given a psychological prompt, generate a step-by-step reasoning process only.
  • Do not output the final answer.
  • Do not include <think>or <answer>tags.
  • If it is a direct emotion prediction, use only a few words.
  • Generate reasoning for [QA] with the expected answer [Answer].
A.2 Other rewards
For our base reward, each group $(b,i)$ is assigned a final reward $r^{\text{Final}}_{b,i}$ as follows:
$r^{\text{Final}}_{b,i}=w^{\text{F1}}\cdot r^{\text{F1}}_{b,i}+w^{\text{fmt}}\cdot r^{\text{fmt}}_{b,i}$ (11)
All components are aligned with the method section above. For the length or short reward, each group $(b,i)$ is assigned a final reward $r^{\text{Final}}_{b,i}$ as follows:
$r^{\text{Final}}_{b,i}=w^{\text{F1}}\cdot r^{\text{F1}}_{b,i}+w^{\text{fmt}}\cdot r^{\text{fmt}}_{b,i}+w^{\text{length/short}}\cdot r^{\text{length/short}}_{b,i}$ (12)
For the length reward, we have the condition:
$r^{\text{length}}_{b,i}\!=\!\begin{cases}\log\left(1+\frac{r^{\text{F1}}_{b,i}}{\bar{r}^{\text{F1}}}\right)\!\!\cdot\!\!\frac{L_{b,i}}{\bar{L}},\!\!\!\!\!\!\!&\frac{L_{b,i}}{\bar{L}}\!\leq\!L_{+}\\ 0,&\text{otherwise}\end{cases}$ (13)
For the short reward, we have the condition:
$r^{\text{short}}_{b,i}\!=\!\begin{cases}\log\left(1+\frac{r^{\text{F1}}_{b,i}}{\bar{r}^{\text{F1}}}\right)\!\!\cdot\!\!\frac{L_{b,i}}{\bar{L}},\!\!\!\!\!\!\!&\frac{L_{b,i}}{\bar{L}}\!\geq\!S_{-}\\ 0,&\text{otherwise}\end{cases}$ (14)
For the length reward with repetition penalty, each group $(b,i)$ is assigned a final reward $r^{\text{Final}}_{b,i}$ as follows:
$r^{\text{Final}}_{b,i}=w^{\text{F1}}\cdot r^{\text{F1}}_{b,i}+w^{\text{fmt}}\cdot r^{\text{fmt}}_{b,i}+w^{\text{length}}\cdot r^{\text{length}}_{b,i}-r^{\text{rep}}_{b,i}$ (15)
All components are consistent with the descriptions above. Beyond the reward functions described above, we designed two alternative bilateral rewards that both adapted logarithmic functions to encourage the reasoning length to stay within a desirable range. However, while these reward methods performed reasonably well on the training set, their performance on the OOD datasets remained limited, as the generated reasoning lengths were overly constrained to the distribution of our own dataset.
A.3 GRPO parameters
For training our models, we used the following parameter set accross for all our GRPO/TGRPO training:
  • Model: Qwen2.5-0.5/1.5/3B-Instruct
  • Data Input:
    • Training/Validation files: train, test
    • Prompt/Answer keys: prompt, answer
    • Max prompt/response length: 2048 tokens
    • Overlong prompts filtered: true
  • Optimization:
    • Optimizer: AdamW with learning rate $1\times 10^{-6}$ and weight decay $1\times 10^{-2}$
    • Gradient clipping: max norm of 1.0
    • Learning rate warmup ratio: 0.0
  • Batching and Parallelism:
    • Global batch size: 128
    • Micro batch size (update): 4 per device
    • Micro batch size (experience): 16 per device
    • Rollout batch size: 512; Validation batch size: 1024
    • Tensor parallel size: 2
    • FSDP: full sharding with rank-0 init enabled; CPU offload enabled for reference model
    • Offloading: parameters and optimizer offloaded to CPU
  • Algorithm: GRPO (Group Relative Policy Optimization)
    • KL regularization enabled with coefficient $1\times 10^{-2}$
    • KL penalty type: low_var_kl
    • Advantage estimator: grpo
  • Rollout Strategy:
    • Samples per prompt: 5
    • Temperature: 1.0 (0.5 for validation)
    • Top-p sampling: 0.99
  • Reward Function: Custom Python-based function at math.py:compute_score
  • Training Configuration:
    • Steps: 315
    • GPU utilization rate: 0.8
A.4 SFT parameters
We fine-tuned our models in the SFT (Supervised Fine-Tuning) stage using the following configuration:
  • Model: Qwen2.5-0.5/1.5/3B-Instruct
  • Fine-Tuning Setup:
    • Finetuning type: full
    • DeepSpeed configuration: ds_z3_config.json
    • Precision: bf16 enabled
    • Training stage: sft
  • Dataset:
    • Dataset name: llama with qwen template
    • Maximum sequence length: 2048 tokens
    • Preprocessing: 16 workers for data prep, 4 for dataloader
  • Training Configuration:
    • Total steps: 315
    • Learning rate: $1\times 10^{-5}$ with cosine scheduler
    • Warmup steps: 31 (10% of total)
    • Batch size: 1 per device, with gradient accumulation steps = 2
    • DDP timeout: 180000000
Ethics and Privacy Statement
To minimize all potential risks of unintentional disclosure, we have implemented strict organizational and technical controls.
First, none of the collected stories were ever shared with anyone beyond those directly involved in data collection and primary research—i.e., no individual or institution outside the designated research team accessed the full stories in any shape or form. Second, all data processing was conducted solely on secure local servers within our institution. We did not use any cloud services or third‑party platforms; no private data has ever been accessed by parties not explicitly authorized. Throughout the entire process, we continuously assessed re‑identification risk, drawing on extensive literature documenting the fragility of “anonymous” data against re‑identification attacks. Whenever any potential residual risk persisted, additional redaction or complete removal was undertaken. Importantly, every step of story collection and annotation was carried out by trained psychology students, under the direct supervision of graduate students or faculty members in psychology, ensuring both methodological rigor and strict adherence to ethical standards.
These organizational and technical safeguards ensure that at no point did anyone outside the designated collection team view identifiable participant stories—reflecting our commitment to privacy, responsibility, and the public good as emphasized by AAAI’s Ethics and Plurality principles.
Appendix B Experimental Details
Data Preparation and Comparison Datasets.
We build our experimental datasets starting from StimuliQA, which contains 3,280 expert-annotated narratives covering 54 psychological variables. To increase topic diversity and mitigate repetition, we manually extract 10–50 candidate topics per variable based on sample size, and then expand them into up to 200 diverse topics each using in-context learning (ICL) prompts with LLaMA 3.3 Instruct 70B. Following the same pipeline, we construct two external comparison datasets: MistralQA (from Mistral 8$\times$7B) and LlamaQA (from LLaMA 3.3 Instruct 70B). These datasets collectively provide a broad coverage of psychological reasoning scenarios for pre-training and evaluation.
Logical Reasoning Sample Injection.
To enhance reasoning ability without altering the output format, we curate a reasoning-focused subset. Specifically, Llama 3.3 Instruct 70B annotates 500 short and 500 long reasoning samples from StimuliQA with explicit thinking tokens. These 1,000 samples are then used to SFT-train Qwen models (0.5B, 1.5B, and 3B), effectively injecting professional psychological reasoning knowledge from StimuliQA while preserving compatibility with downstream evaluation.
Reward Design and Bilateral RL Training.
Our reinforcement learning framework adopts a Bilateral Reward (BR) that integrates correctness (F1-based), format adherence, length-awareness, and repetition penalties. This reward structure encourages the model to produce accurate, well-structured reasoning, adaptively control reasoning length, and avoid redundant outputs. For GRPO training, we select the SFT checkpoint that achieves the median F1 on StimuliQA to ensure stable learning.
Continual Learning Module.
We implement a filtering-based continual learning procedure to simulate user-specific feedback. Predictions from Psy-Interpreter on the validation set are retained only if their score exceeds 0.4; the predicted answer is then replaced with the most F1-aligned choice (or random for “E” cases). These curated samples are used for an additional SFT step to produce dataset-specific models that support incremental adaptation across tasks.
Training Sets ToMbench SimpleTom Social IQA Cosmos QA BIG-Bench Hard
F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Qwen2.5-0.5B-Instruct
Llama3.3-70B 20.85 31.39% 21.13 31.50% 32.49 42.90% 19.75 29.97% 14.62 37.23%
Mistral8$\times$7B 20.51 29.99% 29.05 32.93% 34.02 36.48% 15.93 27.51% 19.90 38.25%
Human Labeled 23.70 35.51% 29.89 41.41% 37.90 51.09% 21.36 33.42% 20.01 38.98%
Qwen2.5-1.5B-Instruct
Llama3.3-70B 21.40 38.66% 21.00 31.56% 39.14 57.10% 19.42 32.65% 19.33 47.74%
Mistral8$\times$7B 23.98 40.17% 33.45 27.93% 42.79 48.70% 17.40 29.54% 19.39 38.10%
Human Labeled 25.88 47.55% 34.51 59.08% 46.42 57.56% 23.24 39.72% 19.60 48.47%
Qwen2.5-3B-Instruct
Llama3.3-70B 25.40 42.13% 51.93 49.20% 53.73 64.56% 20.94 35.07% 16.02 42.77%
Mistral8$\times$7B 25.17 42.58% 54.38 38.27% 58.39 63.37% 17.48 32.00% 15.89 34.60%
Human Labeled 25.55 49.69% 55.31 51.76% 64.28 68.81% 23.06 39.22% 16.14 53.87%
Table 4: Comparison of GRPO R1 training on StimuliQA and two other comparion datasets.
Dataset Expert Psych. Stimuli Real‑world Varied Large Varied General-
Annotation Reasoning Authenticity Prevalence Difficulty Scale Tasks izability
ToMBench yes yes no partial yes no yes yes
SimpleToM partial yes partial yes yes no yes yes
SocialIQa no yes no yes no yes no yes
CosmosQA no yes yes yes no yes partial yes
BIG-Bench Hard yes yes yes no no no yes yes
MHQA partial yes yes yes no yes yes no
Psych-101 yes yes NA yes yes yes yes no
BoltMonkey yes no NA NA yes yes yes yes
MentalChat16K partial yes partial yes yes no no yes
StimuliQA yes yes yes yes yes yes yes yes
Table 5: Comparison of psychology-focused QA datasets across multiple criteria. “Partial” indicates that the dataset partially fulfills the requirement, whereas “NA” denotes that the criterion does not apply to the dataset.
Dataset Construction and Quality Control.
During dataset construction, we observed that LLMs sometimes bypass genuine reasoning by directly leveraging human-labeled psychological parameters. For example, if a Stimuli carries a B_disap score of 2, the model might simply rephrase the parameter into a question and return the annotated value as the answer, achieving a correct response without engaging in meaningful inference. To prevent such shortcut behavior, we systematically filtered out trivial QA pairs and instead adopted chain-of-thought prompting with multi-turn dialogue generation, ensuring that each question requires narrative-grounded reasoning rather than mere label retrieval. This refinement not only supports explicit knowledge injection during training but also enhances the dataset’s ability to drive robust generalization in psychological reasoning tasks, enabling models to perform well in out-of-distribution scenarios while maintaining a diverse spectrum of reasoning difficulty.
Difficulty Assessment.
Before model training, we conducted a comprehensive assessment to determine whether StimuliQA’s intrinsic difficulty distribution could bias reasoning length and model attention over time. To this end, we first categorized a subset of stories based on psychological parameter density and emotional complexity, as these factors directly influence the depth of reasoning required. Next, we prompted a larger‑scale LLM to produce explicit step‑by‑step reasoning sequences for randomly sampled stories, and measured the sequence length by word count as a proxy for cognitive effort.
Across 10 independent sampling rounds, where the decoding temperature was gradually increased from 0.1 to 1.0 to simulate both deterministic and creative reasoning behaviors, the reasoning lengths ranged from 87 to 136 words, with an overall average of 104 words. We also observed that stories containing multi‑layered social interactions, such as overlapping intentions or conflicting emotions, tended to elicit longer reasoning chains, whereas single‑event narratives generated more concise reasoning. Interestingly, the variance across rounds remained relatively small, suggesting that the dataset itself imposes a consistent cognitive demand regardless of the sampling temperature. This implies that the diversity in StimuliQA primarily comes from intrinsic story structure rather than stochastic model generation.
This analysis suggests that StimuliQA exhibits a moderately broad difficulty distribution without extreme skew, making it suitable for length‑sensitive training strategies in reinforcement learning. Moreover, the stable word‑length distribution provides a natural signal for detecting overfitting or reasoning shortcuts during model evaluation, and it ensures that curriculum or reward‑based training can be calibrated without the risk of abrupt difficulty spikes. Together, these findings confirm that StimuliQA provides a balanced environment for modeling psychologically grounded reasoning, enabling both robust training and interpretable evaluation.
Evaluation Protocol.
Psy‑Interpreter is evaluated under a unified protocol across five out‑of‑distribution (OOD) benchmarks covering theory‑of‑mind inference, commonsense reasoning, and emotionally grounded psychological interpretation. Specifically, we include tasks from ToMBench and SimpleToM to assess mental‑state attribution, SocialIQA and CosmosQA to evaluate narrative‑driven commonsense and causal reasoning, and BIG‑Bench Hard to test extreme generalization to low‑frequency psychological scenarios. For each benchmark, we report both accuracy and F1 score: accuracy reflects the model’s ability to select the correct answer in a multiple‑choice or span‑selection setting, while F1 additionally captures partial correctness and robustness to distractor options, which are critical in psychologically nuanced QA. Evaluation is conducted using a standardized inference pipeline with temperature‑controlled decoding to reduce randomness, and all predictions are cross‑checked for format consistency. This protocol ensures that observed performance differences arise from genuine reasoning capability rather than prompt formatting or spurious correlations, thereby providing a reliable measure of Psy‑Interpreter’s generalization across diverse reasoning environments.
Appendix C Dataset and Code
Please check our anonymous repo via the link https://anonymous.4open.science/r/Stimuli2Minds-E37B for our available datasets and code. Our experiment results will be further listed in this given repository.
C.1 Datasets Advantages
In this subsection, we present the advantages of StimuliQA and underscore the importance of building such a dataset. We evaluate the current open-source psychology-related QA datasets alongside ours from eight complementary perspectives:
  1. 1.
    Expert Annotation – Datasets annotated by psychology experts provide high-fidelity supervision and inject human-like interpretive signals that are difficult to achieve with crowd-sourcing or synthetic labels. This ensures that subtle emotional and cognitive cues are correctly captured.
  2. 2.
    Psychological Reasoning – Beyond surface-level comprehension, psychological reasoning measures whether a dataset requires models to infer latent mental states, intentions, and social-cognitive processes. StimuliQA is explicitly designed to elicit such reasoning, enabling models to move beyond pattern matching.
  3. 3.
    Stimuli Authenticity – Authentic narratives, drawn from real or plausibly human experiences, reduce domain gaps between training and real-world usage. They help models learn to interpret natural, context-rich scenarios instead of fabricated or overly simplified prompts.
  4. 4.
    Real-world Prevalence – This criterion reflects whether the events or scenarios in the dataset are representative of common psychological experiences in everyday life. Datasets with high prevalence promote better model generalization to real-world applications.
  5. 5.
    Varied Difficulty – Psychological reasoning is inherently multi-layered, ranging from explicit emotion recognition to complex Theory-of-Mind inference. A dataset that spans multiple difficulty levels facilitates curriculum learning and prevents overfitting to trivial cases.
  6. 6.
    Large Scale – Larger datasets provide statistical coverage across diverse mental states and social situations, which is crucial for training robust models. StimuliQA achieves scale without sacrificing annotation quality.
  7. 7.
    Varied Tasks – Datasets that support multiple task formulations (e.g., multiple-choice, open-ended QA, emotion tagging) foster broader capability and flexible evaluation of psychological reasoning in LLMs.
  8. 8.
    Generalizability – The ultimate measure of a dataset’s value lies in whether models trained on it can transfer to unseen psychological tasks and out-of-distribution (OOD) scenarios. StimuliQA is explicitly constructed to maximize cross-task and cross-domain generalization.
By excelling across all eight dimensions (Table 5), StimuliQA provides comprehensive and theory-grounded coverage that existing datasets lack, serving as a foundational resource for advancing psychologically informed language models. These advantages collectively justify our choice to construct an expert-annotated dataset rather than relying solely on public resources.
Building on this foundation, StimuliQA is deliberately designed to capture a microcosm of society. We collected stimuli spanning multiple age groups to reflect life-stage differences in psychological reasoning, and for each ethical theme, we ensured coverage across genders to avoid demographic bias. Furthermore, our dataset incorporates narratives from both long-term residents and recent immigrants, offering a nuanced representation of cultural integration and social diversity. This multi-dimensional coverage enables StimuliQA not only to support psychological reasoning tasks, but also to serve as a miniature model of societal dynamics, enriching downstream analyses of human-like cognition and socially grounded inference.
Moreover, our decision to convert realistic psychological narratives into QA format is not merely a matter of evaluation convenience. QA pairs compel models to perform explicit reasoning steps that mirror real-world human cognition: interpreting events, inferring latent emotions, and formulating socially aware responses. Unlike raw story corpora or unstructured annotations, QA-formatted stimuli ensure that psychological reasoning can be quantitatively measured, reproducibly tested, and directly applied to practical domains such as mental health triage, social behavior analysis, or human-centric AI assistants. This design bridges the gap between academic research and applied impact, making StimuliQA a high-value resource for real-world, psychologically informed AI systems. In essence, StimuliQA not only trains models to understand humans better but also creates a systematic pathway for deployable, socially responsible cognitive reasoning.
C.2 Datasets Samples
Tables 610 show sample entries from MistralQA and LlamaQA, which serve as examples demonstrating the format of our dataset.
Prompt & Question
Stimuli: I recently attended a local political rally to support my preferred candidate. As the event came to a close, a group of radical protestors, who opposed our beliefs, started causing chaos and disrupting the peace. They began shouting slogans with aggressive undertones, which made me feel uneasy and disappointed. Even though I fundamentally disagreed with their views, I couldn’t help but worry about the escalation of tensions and fear the potential for political extremism.
Q: Which emotion best describes your reaction to the political protestors?
Correct Answer: Disappointment directed towards political entities
Table 6: Example of a prompt and its corresponding correct answer illustrating C_dpol.
Prompt & Question
Stimuli: A few months ago, I participated in a local fundraising event for cancer research. The organizers did a fantastic job raising awareness about early detection methods and the importance of regular check-ups through their well-coordinated public awareness campaign. Their efforts inspired me to spread the word among my friends and family, who were also grateful for the information.
Q: Which aspect of the public awareness campaign evoked feelings of gratitude directed towards social collective?
Correct Answer: The inspiration I felt to inform others about early detection methods.
Table 7: Example of a prompt and its corresponding correct answer illustrating C_gcol.
Prompt & Question
Stimuli: As I walked into my new job at a tech firm, I was excited to join the coding team. On my first day, I noticed that our team lead, Rachel, made a point to introduce me to every member of the team, highlighting their unique backgrounds and specialties. She mentioned that Maria brings a Latin American perspective to her approach to algorithm design, Jake applies his knowledge of African cultural patterns to create more inclusive UI/UX, and Leila utilizes her experience growing up in India to develop innovative solutions for international clients. Rachel emphasized how these diverse viewpoints have been instrumental in helping our company stand out in the industry by creating software that caters to a wide range of users from different ethnic backgrounds. I felt inspired by the emphasis on celebrating and leveraging our differences to drive innovation.
Q: What aspect of the coding team’s dynamics is most reflective of the psychological parameter ’J_div’ (diversity mentioned in the coding unit), as observed in the Stimuli?
Correct Answer: The intentional highlighting and utilization of team members’ diverse ethnic backgrounds and experiences in coding approaches.
Table 8: Prompt reflecting team diversity and its psychological parameter annotation for parameter J_div (diversity mentioned in the coding unit),.
Prompt & Question
Stimuli: I went to the beach yesterday. It was a place with sand and water. I saw people doing things. Some were swimming, others were sunbathing. I walked along the shore for a bit, then left. My day was otherwise uneventful.
Q: How would you rate my emotional engagement or detachment from the scene I described at the beach?
Correct Answer: Completely detached, describing the scene in a purely factual manner without any emotional tone.
Table 9: Prompt illustrating emotional detachment annotation.
Prompt & Question
Stimuli: I had been having an ongoing disagreement with my neighbor over his dog constantly barking and disturbing our peace. Today, he accused me of not taking proper care of my garden, which was far from the truth. I felt a surge of anger as I recalled all the times I had tried to discuss the dog issue with him.
Q: How would you describe the speaker’s emotional state in this scenario?
Correct Answer: Very angry.
Table 10: Prompt reflecting high anger in neighbor’s dog.
C.3 Model reasoning sample
The following presents a comparison of model reasoning on ToMBench’s Unexpected Outcome Test, where Index 1 typically involves more straightforward emotional predictions, while Index 2 and 3 require contrastive or unexpected reasoning, making this benchmark a clear example for analyzing reasoning robustness. To comprehensively compare the reasoning processes of the two reward systems, we selected one representative sample from each index for demonstration, where both the base model and the bilateral model produced correct (or relatively correct) answers. For fairness, neither model received any additional SFT knowledge injection; they were evaluated purely as Qwen/Qwen2.5‑3B‑Instruct after RL training, ensuring that any differences in reasoning can be attributed directly to the reward design rather than prior knowledge advantages.
Stimuli
Mrs. Li’s son, Xiao Qiang, is in college now, he tells her that he is the captain of the school football team. This achievement is high because the football team is very popular in school.
Question
How does Mrs. Li react when Xiao Qiang is chosen as the football team captain?
Base Interpretation
Reasoning: In this story, Mrs. Li’s son, Xiao Qiang, has been chosen as the captain of the school football team. This is significant because the football team is very popular in school, suggesting that Xiao Qiang has demonstrated exceptional skills and leadership. Mrs. Li, as a parent, would likely react with a mix of pride, joy, and perhaps some excitement. She would probably be eager to learn more about his achievements and the challenges he faced in being chosen as the captain. She might also be excited about the opportunities this brings for her son to interact with other students and to potentially develop his leadership skills further.
Bilateral Interpretation
Reasoning: Mrs. Li would likely be very proud of her son for being chosen as the football team captain.
Answers
Ground Truth: Proud
Base Answer: With pride
Bilateral Answer: She is proud of him
Table 11: Comparison between the Base model, Bilateral model, and ground truth on a ToMbench sample (Index1).
As shown in Table 11, the Bilateral Interpretation provides a more accurate and psychologically grounded reasoning compared to the Base model. From a Theory-of-Mind perspective, effective emotional inference requires identifying the primary emotional response of the agent—in this case, Mrs. Li, without being distracted by peripheral or speculative reasoning. The Base model generates a verbose chain of thought, speculating about potential excitement, curiosity, and leadership development opportunities. However, such reasoning, while plausible, introduces extraneous mental states that are not directly supported by the stimulus, diluting the clarity of the predicted reaction.
By contrast, the Bilateral Interpretation adopts a more direct and focal reasoning strategy, immediately linking the salient event, “Xiao Qiang is chosen as the football team captain”, to its core psychological implication for the parent: pride. This aligns with established findings in social and developmental psychology that parental emotional reactions to a child’s social achievements are typically dominated by pride rather than curiosity or anxiety. In other words, the Bilateral model demonstrates a reasoning style that suppresses irrelevant inferences and highlights the dominant emotional appraisal, leading to a prediction that is both concise and aligned with the ground truth.
Stimuli
Mrs. Li’s son, Xiao Qiang, is in college now, he tells her that he is the captain of the school football team. This achievement is high because the football team is very popular in school.
Question
Mrs. Li should be proud of her son becoming the football team captain, but why does she actually seem worried?
Base Interpretation
Reasoning: Mrs. Li might be worried for a few reasons. First, she may be concerned about the pressure Xiao Qiang faces as a captain, including the expectations from his teammates, coaches, and school administration. Second, she might be worried about his safety and the risks associated with playing football, especially if she is concerned about injuries. Additionally, she might be worried about his academic performance, as balancing sports and studies can be challenging. Lastly, she might be concerned about his well-being and the emotional toll of high-pressure situations.
Bilateral Interpretation
Reasoning: Mrs. Li would likely be very proud of her son for being chosen as the football team captain.
Answers
Ground Truth: Her son may face pressure in balancing study and football activities.
Base Answer: Because she is worried about his academic and emotional well-being.
Bilateral Answer: Because she is worried about his safety and the pressure of the role.
Table 12: Comparison between the Base model, Bilateral model, and ground truth on a ToMbench sample (Index2).
As shown in Table 12, both the Base and Bilateral interpretations attempt to explain Mrs. Li’s worry despite her pride in Xiao Qiang’s achievement. The Base model produces a long, enumerative reasoning sequence, speculating on multiple possible concerns—academic performance, physical safety, emotional stress, and social expectations. While this reasoning is broadly plausible, it reflects a diffuse cognitive pattern in psychological terms, where the model distributes attention across many secondary emotional states instead of isolating the dominant source of worry. This over-generation risks introducing irrelevant inferences that dilute alignment with the ground truth.
In contrast, the Bilateral Interpretation adopts a focused and dual-channel reasoning strategy aligned with social-cognitive psychology. It highlights the two most salient drivers of parental worry: role pressure and safety concerns. These correspond to the primary appraisal dimensions in Lazarus’ cognitive-motivational-relational theory of emotion, where parents appraise both the demands of the new social role and the potential risks it entails. By selectively attending to the most psychologically plausible causes of worry, the Bilateral model exhibits higher signal-to-noise ratio in emotional inference and aligns more closely with the ground truth reasoning process.
Stimuli
Mrs. Li’s son, Xiaoqiang, goes to college and tells her that he is the captain of the school football team. This achievement is high because the football team is very popular in school. But Mrs. Li thinks her son may face pressure in balancing study and football activities. Xiaoqiang tells Mrs. Li that he is the captain of the football team.
Question
What emotion does Mrs. Li possibly show?
Base Interpretation
Reasoning: In this story, Mrs. Li hears that her son, Xiaoqiang, is the captain of the school football team. While this is undoubtedly a positive achievement, Mrs. Li might also be concerned about the potential stress and balancing act Xiaoqiang faces between his studies and his football commitments. Given her concern, it’s likely that Mrs. Li feels a mix of pride and worry. She may be concerned about the pressure Xiaoqiang might be under and if he is managing his time effectively.
Bilateral Interpretation
Reasoning: Based on the story, Mrs. Li’s son, Xiaoqiang, tells her that he is the captain of the school football team, which is a significant achievement. However, Mrs. Li might be concerned about the pressure Xiaoqiang may face in balancing his studies and football activities. Given this concern, it’s likely that Mrs. Li is worried or anxious about her son’s situation.
Answers
Ground Truth: Worried.
Base Answer: Concern and worry.
Bilateral Answer: Worried and concerned.
Table 13: Comparison between the Base model, Bilateral model, and ground truth on a ToMbench sample (Index3).
As shown in Table 13, both the Base and Bilateral interpretations provide partially correct emotional inferences regarding Mrs. Li’s reaction. The Base model identifies a mixed emotional state, suggesting that Mrs. Li feels “a mix of pride and worry.” While such reasoning appears plausible in a general parental context, it demonstrates a projection bias: the model infers pride based on the objective achievement (being a football captain) rather than the actual emotional cues present in the stimulus. Psychologically, this reflects an over-reliance on schema-based inference—assuming that any parental recognition of achievement automatically elicits pride—without sufficiently weighting the narrative evidence.
In contrast, the Bilateral Interpretation aligns more closely with the dominant emotional appraisal principle in affective psychology. It correctly isolates worry and concern as Mrs. Li’s primary emotional response, directly grounded in the stimulus: “Mrs. Li thinks her son may face pressure in balancing study and football activities.” By suppressing unsupported inferences about pride, the Bilateral model exhibits a reasoning pattern that prioritizes contextually verifiable emotional states, resulting in an answer (worried and concerned) that fully matches the ground truth.
This comparison highlights that Bilateral reasoning is not only more concise but also demonstrates selective attention to stimulus-relevant cues, which is critical for Theory-of-Mind (ToM) reasoning tasks. It avoids the common pitfall of overgeneralized emotional attribution observed in the Base model, where the model extrapolates from general parental schemas instead of relying on the narrative evidence.
Conclusion.
Across the three representative samples from Unexpected Outcome Test, our analysis demonstrates that the Bilateral model consistently exhibits a more focused and stimulus-grounded reasoning pattern compared to the Base model. While the Base model frequently generates verbose or schema-driven reasoning, introducing extraneous emotional states such as speculative pride or curiosity, the Bilateral model selectively attends to dominant, contextually verifiable emotions, leading to answers that align more closely with the ground truth. This pattern reflects higher signal-to-noise ratio in emotional inference, improved suppression of irrelevant projections, and enhanced Theory-of-Mind reasoning fidelity, underscoring the effectiveness of bilateral reward design in guiding model reasoning.
Appendix D Limitation
While our current approach leverages length‑based reasoning incentives to guide LLMs toward more structured and cognitively transparent reasoning in psychological tasks, it remains limited in granularity. Relying solely on length encourages step‑by‑step reasoning but may not fully capture the qualitative dimensions of psychological inference, such as the emotional salience or contextual accuracy of mental‑state predictions. A promising future direction is to incorporate fine‑grained reward signals—for example, rewarding accurate emotional appraisal, causal consistency, or alignment with psychological theory—so that the LLM receives multi‑dimensional reinforcement rather than a single length‑based signal. This would enable more precise and nuanced guidance in modeling human‑like mental‑state reasoning.
Appendix E Future Work
In future work, we plan to integrate additional psychology‑oriented reward signals into our reinforcement learning framework, moving beyond length‑based incentives to capture richer dimensions of human mental‑state reasoning. Specifically, we aim to explore emotion‑aligned rewards, causal reasoning rewards, and other psychologically grounded metrics that can provide more precise and theory‑consistent guidance during model training. Moreover, we will investigate whether fine‑grained, psychology‑aware reinforcement learning strategies can further enhance model fidelity and robustness across diverse ToM and affective inference tasks. Another key direction is to examine the generalizability of our approach to broader domains of behavioral and cognitive reasoning, conducting controlled experiments to assess its applicability in tasks such as social interaction prediction, moral judgment, and decision‑making under uncertainty.

### v4

@@
 By integrating real-world stimuli with structured reasoning guidance, our approach enables compact models to internalize social-cognitive principles, perform nuanced psychological inference, and support continual self-improvement.
-Comprehensive experiments across multiple benchmarks further demonstrate that our models achieve expert-level interpretive capabilities, exhibiting strong out-of-distribution generalization and robust continual learning across diverse, challenging, and psychologically grounded tasks.
+Experiments across multiple benchmarks demonstrate that our models achieve expert-level interpretive capabilities, exhibiting strong out-of-distribution generalization across diverse psychologically tasks.
+Our code is publicly available at https://github.com/Githubuseryf/Stimuli2Minds.
 1 Introduction
@@
 2025; Feng et al.
-2025).
+2025b).
 Their emerging potential in psychological domains, such as emotion understanding (Kovacevic et al.
@@
 2025).
-Figure 1: A sample ToMbench question presents two psychological stimuli—one focused on direct emotion prediction, the other requiring nuanced reasoning about mental states.
-Figure 2: The figure summarizes key psychological stimuli across our datasets—emotional variables, narrative transformations, and collective factors—comprising 35,084, 24,831, and 36,890 QA instances, each offering insights into human narratives.
+Figure 1: A sample ToMbench question presents two types of psychological stimuli pairs for demonstration.
+Figure 2: The figure summarizes key psychological parameters across our datasets: emotional variables, narrative transformations, and collective factors, comprising 35,084, 24,831, and 36,890 QA instances, each offering insights into human narratives.
 To evaluate LLMs’ capacity for cognitive reasoning, several benchmarks have emerged.
@@
 2019b).
-In moral reasoning, LLMs may align with or even surpass humans in perceived ethical competence (Dillion et al.
+In moral reasoning, LLMs may align with or even surpass humans in perceived ethical competence (Liu et al.
 2025; Huang et al.
@@
 2.2 CoT and Reinforcement Learning
-Chain-of-Thought (CoT) prompting improves deliberate reasoning by encouraging stepwise thinking (Wei et al.
-2022; Kojima et al.
-2022), and is often combined with RL to align outputs with human preferences (Ouyang et al.
+Chain-of-Thought (CoT) prompting improves deliberate reasoning by encouraging stepwise thinking (Luo et al.
+2025a; Shen et al.
+2025), and is often combined with RL to align outputs with human preferences (Ouyang et al.
 2022).
-Process supervision and gradient-level feedback, such as in DeepSeek-R1, further enhance intermediate reasoning quality (Liu and et al.
-2025).
-Frameworks like Reflexion and Tree of Thoughts introduce self-evaluation and planning to strengthen coherence (Shinn et al.
-2023; Yao et al.
-2023).
+Process supervision and gradient-level feedback, such as in DeepSeek-R1, further enhance intermediate reasoning quality (Luo et al.
+2025c).
+Frameworks like Reflexion and Tree of Thoughts introduce self-evaluation and planning to strengthen coherence (Luo et al.
+2025b; Feng et al.
+2025a).
 Despite progress, challenges remain in balancing fluency with rigor and designing scalable feedback for ambiguous reasoning.
@@
 $P_{b,i}=\frac{|y^{\text{pred}}_{b,i}\cap y^{\text{gold}}_{b,i}|}{|y^{\text{pred}}_{b,i}|},\ R_{b,i}=\frac{|y^{\text{pred}}_{b,i}\cap y^{\text{gold}}_{b,i}|}{|y^{\text{gold}}_{b,i}|},$ (5)
-where $y^{\text{pred}}_{b,i}$ $y^{\text{gold}}_{b,i}$ are token sets after normalization, including lowercase conversion, punctuation and space cleanup.
+where $y^{\text{pred}}_{b,i}$ & $y^{\text{gold}}_{b,i}$ are token sets after normalization, including lowercase conversion, punctuation and space cleanup.
 Format Compliance ($r^{\text{fmt}}$)
@@
 Bilateral Reasoning ($r^{\text{BR}}$)
-[…]
+To encourage informative yet concise reasoning, we introduce a bilateral reward term that jointly considers relative reasoning length and answer quality.
+Let $L_{b,i}=|o_{b,i}|$ denote the output length for group $(b,i)$, and define the batch-level averages as:
 $\bar{L}=\frac{1}{B_{c}G}\sum_{b=1}^{B_{c}}\sum_{i=1}^{G}L_{b,i},\quad\bar{r}^{\text{F1}}=\frac{1}{B_{c}G}\sum_{b=1}^{B_{c}}\sum_{i=1}^{G}r^{\text{F1}}_{b,i}.$ (7)
@@
 To enable self-evaluation and self-improvement under low-resource settings, we propose a simple framework for continuous learning.
-The model not only generates predictions but also assesses their quality and improves over time—producing more user-aligned and psychologically appropriate responses.
+The model not only generates predictions but also assesses their quality and improves over time, producing more user-aligned and psychologically appropriate responses.
 We define the learning criterion as:
-[…]
+$\text{self\_train}(x)\iff\text{valid}(x)\land\text{confidence}(x)>\tau,$ (10)
 where $x$ is a model output, $\text{valid}(x)$ verifies format and $\text{confidence}(x)$ is the model’s estimated confidence score.
@@
 Llama 3.3 70B Instruct 27.42 52.24% 76.43 61.46% 60.91 69.07% 15.27 37.19% 17.61 69.34%
-Mistral 8$\times$7B Instruct 12.03 34.46% 12.06 42.52% 12.14 42.23% 7.49 24.66% 5.39 26.42%
+[…]
 Qwen2.5-0.5B-Instruct
@@
 4.2 RQ1: Effect of Human-Labeled Data
-One of the central hypotheses in our work is that human-labeled data, particularly when annotated by trained psychology students, can inject high-quality domain knowledge into compact models during SFT.
-We evaluate this hypothesis by comparing models trained on our StimuliQA dataset against those trained on synthetic datasets generated by two powerful large models: Llama 3.3-70B and Mistral 8$\times$7B.
-All datasets were used to train Qwen2.5-Instruct models of three different sizes (0.5B, 1.5B, 3B) via SFT, and evaluated on five diverse OOD benchmarks: ToMbench, SimpleToM, SocialIQa, CosmosQA, and BIG-Bench Hard.
-Where we have the final results displayed in Table 1 and Figure 4.
-This comparison reveals performance gaps and tests the transferability of psychological reasoning across model sizes and tasks.
-These benchmarks span theory of mind, social commonsense, and logical reasoning, allowing for a comprehensive assessment of domain generalization and model capabilities.
-As shown in Table 1, models trained on StimuliQA consistently outperform those trained on synthetic data across all model sizes and benchmarks, with especially strong gains in the 3B setting.
-For instance, on SimpleToM, Qwen2.5-3B achieves 37.62 F1 and 56.44% accuracy, far exceeding Llama 3.3 (18.48/26.16%) and Mistral 8$\times$7B (35.02/33.33%).
-These substantial margins further reinforce our claim that: human-labeled data conveys more nuanced and structured psychological knowledge than synthetic outputs.
-The consistent OOD improvements indicate this knowledge is indeed generalizable, highlighting the professional quality and domain relevance of our dataset.
-Annotations by trained psychology students provide valuable domain expertise essential for effective model training.
-When used in GRPO training (Figure 4), human-labeled data still offers an advantage.
+[…]
+[…]
 4.3 RQ2: Impact of Reward Design
@@
 Data Preparation and Comparison Datasets.
-We build our experimental datasets starting from StimuliQA, which contains 3,280 expert-annotated narratives covering 54 psychological variables.
-To increase topic diversity and mitigate repetition, we manually extract 10–50 candidate topics per variable based on sample size, and then expand them into up to 200 diverse topics each using in-context learning (ICL) prompts with LLaMA 3.3 Instruct 70B.
-Following the same pipeline, we construct two external comparison datasets: MistralQA (from Mistral 8$\times$7B) and LlamaQA (from LLaMA 3.3 Instruct 70B).
-These datasets collectively provide a broad coverage of psychological reasoning scenarios for pre-training and evaluation.
+[…]
 Logical Reasoning Sample Injection.
@@
 Llama3.3-70B 20.85 31.39% 21.13 31.50% 32.49 42.90% 19.75 29.97% 14.62 37.23%
-Mistral8$\times$7B 20.51 29.99% 29.05 32.93% 34.02 36.48% 15.93 27.51% 19.90 38.25%
+[…]
 Human Labeled 23.70 35.51% 29.89 41.41% 37.90 51.09% 21.36 33.42% 20.01 38.98%
@@
 Llama3.3-70B 21.40 38.66% 21.00 31.56% 39.14 57.10% 19.42 32.65% 19.33 47.74%
-Mistral8$\times$7B 23.98 40.17% 33.45 27.93% 42.79 48.70% 17.40 29.54% 19.39 38.10%
+[…]
 Human Labeled 25.88 47.55% 34.51 59.08% 46.42 57.56% 23.24 39.72% 19.60 48.47%
@@
 Llama3.3-70B 25.40 42.13% 51.93 49.20% 53.73 64.56% 20.94 35.07% 16.02 42.77%
-Mistral8$\times$7B 25.17 42.58% 54.38 38.27% 58.39 63.37% 17.48 32.00% 15.89 34.60%
+[…]
 Human Labeled 25.55 49.69% 55.31 51.76% 64.28 68.81% 23.06 39.22% 16.14 53.87%
@@
 Appendix C Dataset and Code
-Please check our anonymous repo via the link https://anonymous.4open.science/r/Stimuli2Minds-E37B for our available datasets and code.
-Our experiment results will be further listed in this given repository.
+Our experiment results will be further listed in our repository.
 C.1 Datasets Advantages
@@
 Stimuli
-Mrs. Li’s son, Xiao Qiang, is in college now, he tells her that he is the captain of the school football team. This achievement is high because the football team is very popular in school.
+Mrs. Li’s son, Xiao Qiang, is now in college. He tells her that he has become the captain of the school football team, a notable achievement since the team is highly popular on campus.
 Question
-Mrs. Li should be proud of her son becoming the football team captain, but why does she actually seem worried?
+Although Mrs. Li should feel proud of her son becoming the football team captain, why does she seem worried?
 Base Interpretation
 Reasoning: Mrs.
-Li might be worried for a few reasons.
-First, she may be concerned about the pressure Xiao Qiang faces as a captain, including the expectations from his teammates, coaches, and school administration.
-Second, she might be worried about his safety and the risks associated with playing football, especially if she is concerned about injuries.
-Additionally, she might be worried about his academic performance, as balancing sports and studies can be challenging.
-Lastly, she might be concerned about his well-being and the emotional toll of high-pressure situations.
+Li may be worried for several reasons.
+She might be concerned about the pressure Xiao Qiang faces as a team captain, including high expectations from teammates and coaches.
+She may also worry about his physical safety and the risk of injuries, as well as the challenge of balancing football responsibilities with his academic workload.
+Additionally, she could be anxious about the emotional toll that high-pressure situations might bring.
 Bilateral Interpretation
-Reasoning: Mrs. Li would likely be very proud of her son for being chosen as the football team captain.
+Reasoning: Mrs. Li is mainly worried that her son will struggle to balance his studies with his demanding role as team captain, and that the added pressure of leadership may affect his well-being.
 Answers
@@
 Base Answer: Because she is worried about his academic and emotional well-being.
-Bilateral Answer: Because she is worried about his safety and the pressure of the role.
-Table 12: Comparison between the Base model, Bilateral model, and ground truth on a ToMbench sample (Index2).
-As shown in Table 12, both the Base and Bilateral interpretations attempt to explain Mrs.
-Li’s worry despite her pride in Xiao Qiang’s achievement.
-The Base model produces a long, enumerative reasoning sequence, speculating on multiple possible concerns—academic performance, physical safety, emotional stress, and social expectations.
-While this reasoning is broadly plausible, it reflects a diffuse cognitive pattern in psychological terms, where the model distributes attention across many secondary emotional states instead of isolating the dominant source of worry.
-This over-generation risks introducing irrelevant inferences that dilute alignment with the ground truth.
-In contrast, the Bilateral Interpretation adopts a focused and dual-channel reasoning strategy aligned with social-cognitive psychology.
-It highlights the two most salient drivers of parental worry: role pressure and safety concerns.
-These correspond to the primary appraisal dimensions in Lazarus’ cognitive-motivational-relational theory of emotion, where parents appraise both the demands of the new social role and the potential risks it entails.
-By selectively attending to the most psychologically plausible causes of worry, the Bilateral model exhibits higher signal-to-noise ratio in emotional inference and aligns more closely with the ground truth reasoning process.
+Bilateral Answer: Because she is worried about the pressure of leadership and its impact on his studies.
+Table 12: Comparison of Base model, Bilateral model, and ground truth reasoning on a ToMbench sample (Index 2).
+As shown in Table 12, both the Base and Bilateral models correctly identify that Mrs.
+Li’s worry stems from the tension between pride and concern.
+The Base model produces a lengthy, enumerative reasoning chain, speculating about multiple secondary concerns such as injuries, social expectations, and emotional stress.
+While plausible, this diffuse reasoning pattern risks diluting alignment with the ground truth by introducing peripheral inferences.
+By contrast, the Bilateral model exhibits a focused reasoning strategy, isolating the dominant source of worry: the pressure of leadership and its impact on academic balance.
+This reasoning reflects a higher signal-to-noise ratio in emotional inference and aligns more closely with the ground truth, illustrating how bilateral reinforcement guides the model toward psychologically salient appraisals rather than exhaustive speculation.
 Stimuli
@@
 Appendix D Limitation
-While our current approach leverages length‑based reasoning incentives to guide LLMs toward more structured and cognitively transparent reasoning in psychological tasks, it remains limited in granularity.
-Relying solely on length encourages step‑by‑step reasoning but may not fully capture the qualitative dimensions of psychological inference, such as the emotional salience or contextual accuracy of mental‑state predictions.
-A promising future direction is to incorporate fine‑grained reward signals—for example, rewarding accurate emotional appraisal, causal consistency, or alignment with psychological theory—so that the LLM receives multi‑dimensional reinforcement rather than a single length‑based signal.
-This would enable more precise and nuanced guidance in modeling human‑like mental‑state reasoning.
+While our current approach leverages length-based reasoning incentives to guide LLMs toward more structured and cognitively transparent reasoning in psychological tasks, it remains limited in granularity.
+Relying solely on length encourages step-by-step reasoning but may not fully capture the qualitative dimensions of psychological inference, such as emotional salience or contextual accuracy in mental-state predictions.
+A promising direction is to incorporate fine-grained reward signals—for example, rewarding accurate emotional appraisal, causal consistency, or alignment with psychological theory—so that the LLM receives multi-dimensional reinforcement rather than a single length-based signal.
+Integrating such rewards could help the model balance interpretability with empathy, learning not only to elaborate its reasoning but also to understand why certain inferences are more plausible or emotionally resonant.
+Exploring hierarchical reward structures that jointly optimize structural reasoning and affective understanding may lead to a more holistic, human-aligned reasoning process.
 Appendix E Future Work